# Bi-directional Model Cascading with Proxy Confidence

### v1

@@ 4.3. Cascade Models
  • EntBin: This uses only the small model $M_{S}$, via entropy of the final layer ($g_{\text{ent}}$). This baseline is similar to the DTU method of Varshney and Baral (2022).
  • MaxProb: This uses only the small model $M_{S}$, via the maximum probability of the final layer ($g_{\text{maxP}}$). This baseline is the MaxProb method of Varshney and Baral (2022).
  • Rand: This is the baseline of Varshney and Baral (2022) where the output of $M_{S}$ or $M_{L}$ is selected at random.
We also include an “oracle” version of Bi-Dir, where rather than using the proxy forward confidence for the large model $M_{L}$, we use the actual maximum probability (i.e., $g_{\text{maxP}}(Q_{M_{L}})$. (This is a kind of oracle as under the scenario of interest, the cascade would not have access to this.)
Table 1. Deferral AUC scores for methods and baselines across datasets.
@@
Table 2. Deferral AUC scores at .2 deferral rate for methods and baselines across datasets
@@
Table 3. Required deferral rate $r_{d}$ of Bi-Directional method to surpass system accuracy of MaxProb at .2 and .4 deferral rate with estimated total cost reduction in parentheses.
@@
4.4. Evaluation
Our primary evaluation is measuring the effect on system accuracy as a function of rate of deferral, like Jitkrittum et al. (2023). We plot deferral curves for individual method analysis, and provide AUC scores derived from these curves to allow comparison across methods. Note that these AUC scores differ from those obtained from e.g. ROC curves: their absolute values are determined by the accuracies of $M_{S}$ and $M_{L}$, and are mostly of use in relative ranking. As we are particularly interested in lower deferral rates (i.e., where most of the work is done by the small model and deferral is avoided where possible), we also calculate AUC for sub-intervals of deferral rates $r_{d}\in[0,0.2]$ and $r_{d}\in[0,0.4]$. To aid in interpreting and quantifying improvements indicated by AUC, we also provide the rates of deferral for different models at the same level of system accuracy.
In our secondary evaluation, we evaluate how well calibrated the confidence estimation methods are in order to understand the reasons for the system performance of the various methods. In addition to reliability diagrams, like Ulmer et al. (2024), we use Expected Calibration Error (ECE), Smoothed Expected Calibration Error (smECE), Brier score and AUROC as metrics.
5. Results
5.1. Main Results: System Performance
Table 1 presents the AUC scores for the full deferral curves. Our methods provide the strongest deferral performance across all datasets, consistently outperforming naive and informed baselines. Bi-Dir achieves best results on on ARC-Easy, ARC-Challenge and MMLU datasets with BackInt performing marginally better on BoolQ and CSQA. We also observe that BackInt is a clear improvement over MaxProb and other baselines. Performance of the Bi-Dir is very close to the version with oracle forward confidence for BoolQ, ARC-Easy and ARC-Challenge, indicating that the relatively straightforward proxy model is sufficient there. On the other hand, a more accurate estimate of the large model should improve deferral on MMLU and CSQA, where $M_{L}$ is not very accurate; further refinement of forward confidence modeling should lead to improved deferral on these tasks. (For further context, we include a discussion of performance with respect to a full oracle with perfect knowledge in App D.)
@@ 5.1. Main Results: System Performance
As an illustration of the deferral curves from which the AUC scores are derived, Figure 2 plots cascade system performance against deferral rate using all deferral methods for MMLU. All methods and baselines provide better accuracy gains than chance over the range of delegation rates. The Bi-Dir method outperforms the other methods most in the earlier sections of the curve, avoiding the wasteful deferral of samples likely to be incorrect in both models. This profile is desirable in a typical cascading scenario where the goal is to maximize gains while utilizing the large model sparingly. Deferral curves for all datasets are included in App B.
Table 2 presents the AUC scores for the first 20% of the deferral curves, quantitatively capturing the observation on Figure 2 about important early deferral rates. Bi-Dir is the best method on all non-binary tasks with BackInt also consistently outperforming standard baselines. The large gaps between individual small model and large model performance in some datasets means large model performance parity is only obtained at high deferral rates; however, comparison to baselines and oracle deferral show good deferral decisions particularly at low rates for BackInt and Bi-Dir. Tasks where small model performance is high—BoolQ, ARC(E)—experience the largest gains above Rand in both our methods and informed baselines. AUC scores for the first 40% of the deferral curves show the same pattern (Table 4 in App A).
@@
Figure 2. Deferral curve for MMLU showing performance of each method over all delegation rates.
5.2. Model Calibration
In this section we explore why BackInt and Bi-Dir produce better results overall by looking at model calibration. Figure 3 presents reliability diagrams (Guo et al., 2017) for BackInt calibration alongside MaxProb for baseline comparison. A key advantage of the BackInt method is that it provides a greater spread of samples across confidence levels, facilitating effective deferral decisions rather than conservatively estimating confidence near the mean accuracy. Specifically in Figure 3, BackInt considers a much higher proportion of samples to be high confidence compared to the MaxProb, and those samples are found to be mostly accurate. This pattern is observed in all datasets to varying degrees (reliability diagrams for remaining datasets appear in App C). For an aggregate view, Table 5 in App C presents calibration metrics for the LLaMA-2 13 billion model. Our BackInt method is the best calibrated for all datasets and metrics except for ECE, which is defined similarly to EntBin calibration scores. These metrics quantify the degree to which confidence is aligned with correctness but do not measure the ability to represent a wider range of calibrated confidence values.
Table 6 in App C shows calibration metrics for the LLaMA-2 70 billion model comparing observed targets to the auxiliary model forward confidence estimation. Auxiliary model confidence is poorly calibrated in general, providing only a rough indicator of expected correctness when compared to observed targets. This rough estimate, however, is available prior to use of the large model and is therefore useful in the deferral process, as evidenced in Table 1. We note here that the large and diverse MMLU dataset, which includes questions of varying difficulty and from a wide range of topics can be much more reliably predicted, reaching a correctness AUC of .724 without outputs. Auxiliary model performance as a calibrator is weakest on the BoolQ dataset; we discuss this in App C.
Figure 3. Confidence Histograms and Reliability Diagrams for baseline max probability confidence and internal representations method on the ARC-Easy dataset.
5.3. Ablations and Further Analyses
@@
In addition to using mean bin accuracy as a target for the auxiliary forward model, we investigate using an auxiliary regression model to directly predict individual output probability from the large model and find the difference to be trivial for both calibration and deferral. The binning method yields slightly better calibration metrics and is the chosen method for all the main experiments. We provide metrics for both calibration and system performance in Table 7 in App E.
@@
We examine the sensitivity of the auxiliary forward model to training set size. We conduct the same experiments as above using smaller subsets of the dataset to train the auxiliary model and compare the final deferral performance. While deferral with access to the larger model output probabilities (Oracle Aux.) performs significantly better than the predicted probabilities at all dataset sizes, we find only small degradations in deferral performance using smaller training set sizes for MMLU down to 25% of samples. On this diverse and challenging dataset, a small amount of data can provide a useful rough approximation to forward confidence. Table 10 in App I gives full data.
@@
Length of Delegated Items
The size of deferred items will also have an impact on cost: longer items deferred to the larger model will be more costly. App H gives an analysis of length of deferred items: Bi-Dir delegates shorter items on average. This is likely to be because the deferral model has some knowledge of which items the larger model is uncertain on, which are more likely to be longer.
6. Conclusion
We have shown that a model cascade that bases its deferral decisions on richer notions of confidence derived from internal states of smaller models and from proxy predictions of confidence from larger models can outperform those based on the widely used maximum probability-based confidence. The improvements correspond to reductions in number of items deferred to the large model, and hence cost, by notable margins, up to 42.5% on the datasets explored.
Our current work only looked at models from models from the same family; given observations from Lebovitz et al. (2023) regarding the usefulness of cascade diversity, future work could explore relating confidence of different model types. Other future work could involve extending the classification tasks here to generation, which have their own specific confidence-related complexities (Gupta et al., 2024).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. As noted, we hope it may help in working towards positive societal consequences such as reduced energy consumption in use of LLMs and in wider accessibility to users as a result of lower cost barriers. We do not foresee negative societal consequences.
There are some limitations to our work, in terms of using models from the same series (as opposed to heterogeneous ones), and using one-shot models. We discuss these in App K.
@@ Appendix A Deferral AUC Scores
Table 4. Deferral AUC scores at .4 deferral rate for methods and baselines across datasets
@@ Appendix B Deferral Curves
The deferral curves in Figures 4, 5, 6, and 7 show system performance at all deferral rates for all methods. BackInt and Bi-Dir provide the best accuracy-deferral trade-off in all datasets, particularly through the early sections of each curve.
Figure 4. Deferral curve for ARC-Challenge showing performance of each method over all delegation rates.
Figure 5. Deferral curve for BoolQ showing system performance over delegation rates.
Figure 6. Deferral curve for ARC-Easy showing system performance over delegation rates.
Figure 7. Deferral curve for CSQA showing system performance over delegation rates.
@@ Appendix C Calibration
Table 5. Evaluation of Calibration Methods for LLaMA-2 13b model across all datasets.
@@
Figure 8. Confidence Histograms and Reliability Diagrams for baseline max probability confidence and internal representations method on the ARC-Challenge dataset.
Table 6. Evaluation of Calibration Methods for LLaMA-2 70b model on the ARC-Easy Dataset
@@
Figure 9. Reliability diagrams for backwards internal confidence in LLaMA-2 13b on BoolQ.
Figure 10. Reliability diagrams for backwards internal confidence in LLaMA-2 13b on MMLU.
Figure 11. Reliability diagrams for backwards internal confidence in LLaMA-2 13b on CSQA.
@@ Appendix D Oracle Deferral
While the cascading approach offers significant performance improvements in many contexts, evaluating the quality of the cascading process is challenging. Total system performance savings, area under the deferral curve, cost-savings at performance parity and other metrics in this domain are sensitive to the individual and relative performance of the models in the cascade on the chosen dataset. Since 100% task accuracy is often mathematically impossible given the selected models, one method we find helpful in contextualizing deferral performance changes consideration of the ”Oracle Deferral Curve”, which we define as the maximum possible task performance at any given deferral rate between two models, without changing the outputs of those models. This perfect oracle initially defers only samples that are correct in the large model and incorrect in the small model until there are no more such samples, then defers samples where the models agree, and finally defers losses until 100% of samples are deferred. While this quality of deferral is not itself achievable, it represents an upper bound on the deferral process at every rate. fig. 12 shows the deferral performance of all methods bounded by the perfect deferral curve for the ARC-Challenge dataset.
Figure 12. Deferral curves of examined methods on ARC-Challenge contextualized with perfect deferral.
@@ Appendix E Auxiliary Forward Model Variant
Table 7. Auxiliary forward model performance for bin accuracy vs individual probability outputs on MMLU with LLaMA-2 70b
@@ Appendix F Auxiliary Model Judgment Examples
We present examples of forward confidence estimation in table 8. Examples 1 and 2 belong to the easiest subset of samples. In both cases there is a strong link between the question and the correct answer. The Sun is the primary source of heat in many contexts and thirst is linked to consuming liquids in its definition. Quantitative questions such as Example 3 are considered more difficult, aligning with prior work applying transformer models to simple math questions (Cobbe et al., 2021). Example 4 is considered a sample with very low expected confidence, the correct answer in this case is subtly distinguished from other options that are the subject of ongoing research.
Table 8. Example questions and their estimated difficulty according to auxiliary forward confidence model.
@@ Appendix G Calibration for Llama 7b
In the case of LLaMA-2 7b for ARC-Challenge and other difficult tasks, confidence reflects generally near-chance accuracy for a majority of samples, as indicated in fig. 13. This near chance estimate, while appropriate given the observed performance, limits the viability of intelligent deferral and weakens cascade performance for both baselines and our methods as shown in fig. 14.
Figure 13. Calibration performance for LLaMA-2 7b on ARC-Challenge.
Figure 14. Deferral curves for cascading with LLaMA-2 7b as $M_{S}$ on ARC-Challenge.
Appendix H Delegated Prompt Length
While average prompt lengths are used for estimating relative cost reductions of methods, a more detailed analysis of which samples are delegated under different methods revealed that the baseline method delegates longer samples to the larger model in most cases compared to our method. This means that relative cost reductions are in fact larger than estimated, since computational cost is a function of prompt length and deferral of larger prompts is more costly. Table  12 presents the mean prompt length (in tokens) for samples sent to the larger model using the Bi-Dir vs MaxProb at a .2 deferral rate. The deferral based solely on confidence of the small model may prioritize deferral of long, complex and difficult samples to the large model leading to increased cost per deferred sample. While deferring difficult samples is generally a good idea, a more nuanced approach to deferral can prioritize deferral of samples that are likely to be correct in the large model and avoid wasteful deferral of complex samples that both models will predict incorrectly.
Table 9. Mean prompt length at .2 deferral rate for MaxProb and Bi-directional deferral and average over all datasets.

### v2

@@ 4.3 Cascade
  • EntBin: This uses only the small model $M_{S}$, via entropy of the final layer ($g_{\text{ent}}$). This baseline is similar to the DTU method of Varshney and Baral [48].
  • MaxProb: This uses only the small model $M_{S}$, via the maximum probability of the final layer ($g_{\text{maxP}}$). This baseline is the MaxProb method of Varshney and Baral [48].
  • Rand: This is the baseline of Varshney and Baral [48] where the output of $M_{S}$ or $M_{L}$ is selected at random.
We also include an “oracle” version of Bi-Dir, where rather than using the proxy forward confidence for the large model $M_{L}$, we use the actual maximum probability (i.e., $g_{\text{maxP}}(Q_{M_{L}(x)})$. (This is a kind of oracle as under the scenario of interest, the cascade would not have access to this in a real-world setting as it requires large model outputs at inference time.)
Table 1: Deferral AUC scores for methods and baselines across datasets.
@@
Table 2: Deferral AUC scores at .2 deferral rate for methods and baselines across datasets
@@
Table 3: Required deferral rate $r_{d}$ of Bi-Directional method to surpass system accuracy of MaxProb at .2 and .4 deferral rate with estimated total cost reduction in parentheses.
@@
4.4 Evaluation
Our primary evaluation is measuring the effect on system accuracy as a function of rate of deferral, like Jitkrittum et al. [21]. We plot deferral curves for individual method analysis, and provide AUC scores derived from these curves to allow comparison across methods. Note that these AUC scores differ from those obtained from e.g. ROC curves: their absolute values are determined by the accuracies of $M_{S}$ and $M_{L}$, and are mostly of use in relative ranking. As we are particularly interested in lower deferral rates (i.e., where most of the work is done by the small model and deferral is avoided where possible), we also calculate AUC for sub-intervals of deferral rates $r_{d}\in[0,0.2]$ and $r_{d}\in[0,0.4]$. To aid in interpreting and quantifying improvements indicated by AUC, we also provide the rates of deferral for different models at the same level of system accuracy.
In our secondary evaluation, we evaluate how well calibrated the confidence estimation methods are in order to understand the reasons for the system performance of the various methods. In addition to reliability diagrams, like Ulmer et al. [47], we use Expected Calibration Error (ECE), Smoothed Expected Calibration Error (smECE), Brier score and AUROC as metrics.
5 Results
5.1 Main Results: System Performance
Table 1 presents the AUC scores for the full deferral curves. Our methods provide the strongest deferral performance across all datasets, consistently outperforming naive and informed baselines. Bi-Dir achieves best results on on ARC-Easy, ARC-Challenge and MMLU datasets with BackInt performing marginally better on BoolQ and CSQA. We also observe that BackInt is a clear improvement over MaxProb and other baselines. Performance of the Bi-Dir is very close to the version with oracle forward confidence for BoolQ, ARC-Easy and ARC-Challenge, indicating that the relatively straightforward proxy model is sufficient there. On the other hand, a more accurate estimate of the large model should improve deferral on MMLU and CSQA, where $M_{L}$ is not very accurate; further refinement of forward confidence modeling should lead to improved deferral on these tasks. (For further context, we include a discussion of performance with respect to a full oracle with perfect knowledge in supplementary material  D[52].)
@@ 5.1 Main Results: System Performance
As an illustration of the deferral curves from which the AUC scores are derived, Figure 2 plots cascade system performance against deferral rate using all deferral methods for MMLU. All methods and baselines provide better accuracy gains than chance over the range of delegation rates. The Bi-Dir method outperforms the other methods most in the earlier sections of the curve, avoiding the wasteful deferral of samples likely to be incorrect in both models. This profile is desirable in a typical cascading scenario where the goal is to maximize gains while utilizing the large model sparingly. Figure 3 shows deferral performance on ARC-Easy, in this case the small model is more capable and the large model correctly classifies almost every sample. This scenario makes BackInt based deferral more competitive with Bi-Dir. While BackInt is effective for tasks that are relatively easy for component models, the Bi-Dir method provides efficient deferral for both easy and challenging tasks. Deferral curves for all datasets are included in supplementary material B[52].
Figure 2: Deferral curve for MMLU showing performance of each method over all delegation rates.
Figure 3: Deferral curve for ARC-Easy showing performance of each method over delegation rates.
Table 2 presents the AUC scores for the first 20% of the deferral curves, quantitatively capturing the observation on Figure 2 about important early deferral rates. Bi-Dir is the best method on all non-binary tasks with BackInt also consistently outperforming standard baselines. The large gaps between individual small model and large model performance in some datasets means large model performance parity is only obtained at high deferral rates; however, comparison to baselines and oracle deferral show good deferral decisions particularly at low rates for BackInt and Bi-Dir. Tasks where small model performance is high—BoolQ, ARC(E)—experience the largest gains above Rand in both our methods and informed baselines. AUC scores for the first 40% of the deferral curves show the same pattern (Table 5 in supplementary material A[52]).
To demonstrate consistency we conducted a non-parametric statistical test of significance using bootstrap resampling and report improvements over the strong MaxProb baseline. Using 1000 test-set resamples with replacement, 95% CIs indicate significant AUC improvement over baseline (intervals do not include zero) at all operating points for all datasets using the Bi-Dir method. The BackInt method alone provides statistically significant improvement at almost every operating point with full results in supplementary material I[52].
@@
5.2 Model Calibration
In this section we explore why BackInt and Bi-Dir produce better results overall by looking at model calibration. Figure 4 presents reliability diagrams [13] for BackInt calibration alongside MaxProb for baseline comparison. A key advantage of the BackInt method is that it provides a greater spread of samples across confidence levels, facilitating effective deferral decisions rather than conservatively estimating confidence near the mean accuracy. Specifically in Figure 4, BackInt considers a much higher proportion of samples to be high confidence compared to the MaxProb, and those samples are found to be mostly accurate. This pattern is observed in all datasets to varying degrees (reliability diagrams for remaining datasets appear in supplementary material C[52]).
Figure 4: Confidence Histograms and Reliability Diagrams for baseline max probability confidence and internal representations method on the ARC-Easy dataset.
For an aggregate view, Table 6 in supplementary C[52] presents calibration metrics for the LLaMA-2 13 billion model. Our BackInt method is the best calibrated for all datasets and metrics except for ECE, which is defined similarly to EntBin calibration scores. These metrics quantify the degree to which confidence is aligned with correctness but do not measure the ability to represent a wider range of calibrated confidence values.
Table 7 in supplementary C[52] shows calibration metrics for the LLaMA-2 70 billion model comparing observed targets to the auxiliary model forward confidence estimation. Auxiliary model confidence is poorly calibrated in general, providing only a rough indicator of expected correctness when compared to observed targets. This rough estimate, however, is available prior to use of the large model and is therefore useful in the deferral process, as evidenced in Table 1. We note here that the large and diverse MMLU dataset, which includes questions of varying difficulty and from a wide range of topics can be much more reliably predicted, reaching a correctness AUC of .724 without outputs. Auxiliary model performance as a calibrator is weakest on the BoolQ dataset; we discuss this in supplementary material C[52].
5.3 Ablations and Further Analyses
@@
In addition to using mean bin accuracy as a target for the auxiliary forward model, we investigate using an auxiliary regression model to directly predict individual output probability from the large model and find the difference to be trivial for both calibration and deferral. The binning method yields slightly better calibration metrics and is the chosen method for all the main experiments. We provide metrics for both calibration and system performance in Table 8 in supplementary material E[52].
@@
We examine the sensitivity of the auxiliary forward model to training set size. We conduct the same experiments as above using smaller subsets of the dataset to train the auxiliary model and compare the final deferral performance. While deferral with access to the larger model output probabilities (Oracle Aux.) performs significantly better than the predicted probabilities at all dataset sizes, we find only small degradations in deferral performance using smaller training set sizes for MMLU down to 25% of samples. On this diverse and challenging dataset, a small amount of data can provide a useful rough approximation to forward confidence. Table 4 shows forward model data sensitivity for Bi-Dir on MMLU where forward confidence was most impactful.
Table 4: Bi-directional deferral performance using smaller training sets on MMLU.
AUC (.2) AUC (.4) AUC (full)
Random 0.094603 0.194561 0.528991
Bi-Dir (Aux 25%) 0.097455 0.203643 0.552574
Bi-Dir (Aux 50%) 0.097741 0.204144 0.552880
Bi-Dir (full) 0.097874 0.204550 0.553650
Bi-Dir (Oracle Aux) 0.099757 0.209051 0.558934
@@
Length of Deferred Items
The size of deferred items will also have an impact on cost: longer items deferred to the larger model will be more costly. Supplement H[52] gives an analysis of length of deferred items: Bi-Dir delegates shorter items on average. This is likely to be because the deferral model has some knowledge of which items the larger model is uncertain on, which are more likely to be longer.
6 Conclusion
We have shown that a model cascade that bases its deferral decisions on richer notions of confidence derived from internal states of smaller models and from proxy predictions of confidence from larger models can outperform those based on the widely used maximum probability-based confidence. Our method provides consistent improvement in deferral quality over strong baselines in both easy and difficult benchmark tasks. These improvements correspond to reductions in number of items deferred to the large model for the same overall performance, hence reducing inference costs by notable margins, up to 42.5% on the datasets explored.
Our current work only looked at models from models from the same family; given observations from Lebovitz et al. [24] regarding the usefulness of cascade diversity, future work could explore relating confidence of different model types. Other future work could involve extending the classification tasks here to generation, which have their own specific confidence-related complexities [14].
@@ Appendix A Deferral AUC Scores
Table 5: Deferral AUC scores at .4 deferral rate for methods and baselines across datasets
@@ Appendix B Deferral Curves
The deferral curves in Figures 5, 6, and 7 show system performance at all deferral rates for all methods. BackInt and Bi-Dir provide the best accuracy-deferral trade-off in all datasets, particularly through the early sections of each curve.
Figure 5: Deferral curve for ARC-Challenge showing performance of each method over all delegation rates.
Figure 6: Deferral curve for BoolQ showing system performance over delegation rates.
Figure 7: Deferral curve for CSQA showing system performance over delegation rates.
@@ Appendix C Calibration
Table 6: Evaluation of Calibration Methods for LLaMA-2 13b model across all datasets.
@@
Figure 8: Confidence Histograms and Reliability Diagrams for baseline max probability confidence and internal representations method on the ARC-Challenge dataset.
Table 7: Evaluation of Calibration Methods for LLaMA-2 70b model on the ARC-Easy Dataset
@@
Figure 9: Reliability diagrams for backwards internal confidence in LLaMA-2 13b on BoolQ.
Figure 10: Reliability diagrams for backwards internal confidence in LLaMA-2 13b on MMLU.
Figure 11: Reliability diagrams for backwards internal confidence in LLaMA-2 13b on CSQA.
@@ Appendix D Oracle Deferral
While the cascading approach offers significant performance improvements in many contexts, evaluating the quality of the cascading process is challenging. Total system performance savings, area under the deferral curve, cost-savings at performance parity and other metrics in this domain are sensitive to the individual and relative performance of the models in the cascade on the chosen dataset. Since 100% task accuracy is often mathematically impossible given the selected models, one method we find helpful in contextualizing deferral performance changes consideration of the "Oracle Deferral Curve", which we define as the maximum possible task performance at any given deferral rate between two models, without changing the outputs of those models. This perfect oracle initially defers only samples that are correct in the large model and incorrect in the small model until there are no more such samples, then defers samples where the models agree, and finally defers losses until 100% of samples are deferred. While this quality of deferral is not itself achievable, it represents an upper bound on the deferral process at every rate. fig.˜12 shows the deferral performance of all methods bounded by the perfect deferral curve for the ARC-Challenge dataset.
Figure 12: Deferral curves of examined methods on ARC-Challenge contextualized with perfect deferral.
@@ Appendix E Auxiliary Forward Model Variant
Table 8: Auxiliary forward model performance for bin accuracy vs individual probability outputs on MMLU with LLaMA-2 70b
@@ Appendix F Auxiliary Model Judgment Examples
We present examples of forward confidence estimation in table˜9. Examples 1 and 2 belong to the easiest subset of samples. In both cases there is a strong link between the question and the correct answer. The Sun is the primary source of heat in many contexts and thirst is linked to consuming liquids in its definition. Quantitative questions such as Example 3 are considered more difficult, aligning with prior work applying transformer models to simple math questions [11]. Example 4 is considered a sample with very low expected confidence, the correct answer in this case is subtly distinguished from other options that are the subject of ongoing research.
Table 9: Example questions and their estimated difficulty according to auxiliary forward confidence model.
@@ Appendix G Calibration for Llama 7b
In the case of LLaMA-2 7b for ARC-Challenge and other difficult tasks, confidence reflects generally near-chance accuracy for a majority of samples, as indicated in fig.˜13. This near chance estimate, while appropriate given the observed performance, limits the viability of intelligent deferral and weakens cascade performance for both baselines and our methods as shown in fig.˜14.
Figure 13: Calibration performance for LLaMA-2 7b on ARC-Challenge.
Figure 14: Deferral curves for cascading with LLaMA-2 7b as $M_{S}$ on ARC-Challenge.
Appendix H Deferred Prompt Length
While average prompt lengths are used for estimating relative cost reductions of methods, a more detailed analysis of which samples are deferred under different methods revealed that the baseline method delegates longer samples to the larger model in most cases compared to our method. This means that relative cost reductions are in fact larger than estimated, since computational cost is a function of prompt length and deferral of larger prompts is more costly. Table  10 presents the mean prompt length (in tokens) for samples sent to the larger model using the Bi-Dir vs MaxProb at a .2 deferral rate. The deferral based solely on confidence of the small model may prioritize deferral of long, complex and difficult samples to the large model leading to increased cost per deferred sample. While deferring difficult samples is generally a good idea, a more nuanced approach to deferral can prioritize deferral of samples that are likely to be correct in the large model and avoid wasteful deferral of complex samples that both models will predict incorrectly.
Table 10: Mean prompt length at .2 deferral rate for MaxProb and Bi-directional deferral and average over all datasets.